A company just paid a dividend of $1.70 per share. You expect the dividend to grow 13% over the next year and 9% two years from now. After two years, you have estimated that the dividend will continue to grow indefinitely at the rate of 4% per year. If the required rate of return is 12% per year, what would be a fair price for this stock today? (Answer to the nearest penny.)
Answer #1

D1=(1.7*1.13)=1.921

D2=(1.921*1.09)=2.09389

Value after year 2=(D2*Growth rate)/(Required return-Growth rate)

=(2.09389*1.04)/(0.12-0.04)

=27.22057

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.921/1.12+2.09389/1.12^2+27.22057/1.12^2

=$25.08(Approx).